OpenStudy (anonymous):

Two proofreaders are reading the same work. One finds 50 mistakes, the other 63 mistakes. If 40 of the mistakes were found by both readers, how many mistakes were found.

OpenStudy (anonymous):

A=50 mistakes. B=63 mistakes. A intersection B = 40 mistakes A union B= how many mistakes. So, A union B= A + B - A intersection B A union B=50 + 63 -40 A union B= 50+23 A union B=73 mistakes were there in all.

OpenStudy (anonymous):